# Bridging the Gap between Empirical Welfare Maximization and Conditional Average Treatment Effect Estimation in Policy Learning

### v1

@@ Abstract
The goal of policy learning is to train a policy function that recommends a treatment given covariates to maximize population welfare. There are two major approaches in policy learning: the empirical welfare maximization (EWM) approach and the plug-in approach. The EWM approach is analogous to a classification problem, where one first builds an estimator of the population welfare, which is a functional of policy functions, and then trains a policy by maximizing the estimated welfare. In contrast, the plug-in approach is based on regression, where one first estimates the conditional average treatment effect (CATE) and then recommends the treatment with the highest estimated outcome. This study bridges the gap between the two approaches by showing that both are based on essentially the same optimization problem. In particular, we prove an exact equivalence between EWM and least squares over a reparameterization of the policy class. As a consequence, the two approaches are interchangeable in several respects and share the same theoretical guarantees under common conditions. Leveraging this equivalence, we propose a novel regularization method for policy learning. Our findings yield a convex and computationally efficient training procedure that avoids the NP-hard combinatorial step typically required in EWM.
@@ 1 Introduction
    Equivalence. We establish that EWM over a policy class $\Pi$ is equivalent to least squares over the class ${\mathcal{G}}_{\Pi}\coloneqq\{g=2\pi-1:\pi\in\Pi\}$ for the target $Y_{1}-Y_{0}$. At the empirical level, EWM with an inverse-probability-weighted welfare estimator (or doubly robust (DR) welfare estimator) equals least squares of an IPW (or DR) pseudo-outcome on $g\in{\mathcal{G}}_{\Pi}$.
@@
    Computation. The equivalence yields a convex training objective and allows us to avoid solving the NP-hard problem that arises from maximizing a combinatorial welfare objective.
@@ 2 Setup
Let $e(x)=P(D=1\mid X=x)$ be the propensity score and $m_{d}(x)=E[Y_{d}\mid X=x]$ the conditional mean outcomes. We assume that there exists a constant $0<\epsilon<1/2$ independent of $n$ and that the variables $X$ and $Y$ are bounded.
@@ 3.1 EWM Policy
$\widehat{W}_{n}^{\mathrm{IPW}}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{Y_{i}D_{i}}{e(X_{i})}\pi(X_{i})+\frac{Y_{i}(1-D_{i})}{1-e(X_{i})}\big(1-\pi(X_{i})\big)\right).$
@@
$\widehat{W}_{n}^{\mathrm{DR}}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left(m_{1}(X_{i})\pi(X_{i})+m_{0}(X_{i})\bigl(1-\pi(X_{i})\bigr)+\frac{D_{i}-e(X_{i})}{e(X_{i})(1-e(X_{i}))}\bigl(Y_{i}-m_{D_{i}}(X_{i})\bigr)\bigl(\pi(X_{i})-\tfrac{1}{2}\bigr)\right),$
where $\widehat{m}_{d}$ ($d\in\{1,0\}$) is an estimator of the conditional expected outcome $m_{d}$. If the estimators of the nuisance parameters satisfy the Donsker condition or are constructed via sample splitting (Klaassen, 1987), also called cross-fitting (Chernozhukov et al., 2018), and meet mild convergence rate conditions, then the bias from estimation error vanishes at a faster rate than $1/\sqrt{n}$. We can estimate the bias-correction weight by using Riesz regression or density-ratio estimation, as proposed in Chernozhukov et al. (2024, 2022); Kato (2025a, b).
@@ 3.2 Plug-in Policy
$\widehat{m}_{d}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=d]\big(Y_{i}-f(X)\big),$
where $\mathcal{F}$} is a class of regression functions $f\colon{\mathcal{X}}\to{\mathbb{R}}$, and construct $\widehat{\tau}$ as $\widehat{\tau}=m_{1}-m_{0}$. As another example, we can estimate $\tau(X)$ directly by regressing $X_{i}$ on a variable that is (asymptotically) unbiased for $Y_{1,i}-Y_{0,i}$. For instance, using an IPW estimator for the target variable, we estimate $\tau$ as
$\widehat{\tau}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\sum^{n}_{i=1}\left(\frac{Y_{i}D_{i}}{e(X_{i})}-\frac{Y_{i}(1-D_{i})}{1-e(X_{i})}-f(X)\right),$
@@ 4 Equivalence between EWM and least squares
Given any $\mathcal{G}$}, we have
$\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[\Big(\big(Y_{1}-Y_{0}\big)-g(X)\Big)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[\big(Y_{1}-Y_{0}\big)^{2}-2g(X)\big(Y_{1}-Y_{0}\big)+g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+1\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)\right].$
@@
Continuing,
$\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)+2\big(Y_{1}-Y_{0}\big)\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)/2+(Y_{1}-Y_{0})\right]$
$=\operatorname*{arg\,min}_{\pi\in\left\{\pi(\cdot)=\big(g(\cdot)+1\big)/2\colon g\in{\mathcal{G}}\right\}}E_{P}\left[-\pi(X)Y_{1}-(1-\pi(X))Y_{0}\right].$
We added and subtracted terms that are irrelevant to the optimization.
Finally, by defining ${\mathcal{G}}={\mathcal{G}}_{\Pi}$, the proof is complete. ∎
@@
We can similarly show the empirical version of this equivalence. Recall that the EWM approach trains a policy $\widehat{\pi}_{n}$ as
@@ 4.2 $[0,1]$-policy
We next consider the case where $\pi$ is a $[0,1]$-valued function; that is, $\pi\colon{\mathcal{X}}\to[0,1]$. Let $\widetilde{\Pi}$ be a policy class that contains such policies, $\pi\colon{\mathcal{X}}\to[0,1]$.
@@
In this case, while the optimal policy in the EWM approach is given by $\pi^{*}=\operatorname*{arg\,max}_{\pi\in\widetilde{\Pi}}W(\pi)$, the optimal regressor is $g^{*}=E_{P}[Y_{1}-Y_{0}\mid X]$. Note that $\pi^{*}(X)$ will take the value $1$ or 0 (even if $\Pi$ does not include the first-best policy). We consider the equivalence between the following two problems: regularized EWM and modified least squares.
We define the regularized EWM as follows:
@@
$\widetilde{g}(\lambda)=2\widetilde{\pi}(4\lambda)-1.$
@@
which also implies $\widetilde{g}(\lambda)\to 2\pi^{*}-1$ as $\lambda\to\infty$.
@@
For any $\mathcal{G}$}, we have
$\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[\left(\frac{1}{\sqrt{\lambda}}\big(Y_{1}-Y_{0}\big)-\sqrt{\lambda}g(X)\right)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[\frac{1}{\lambda}\big(Y_{1}-Y_{0}\big)^{2}-2g(X)\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}E_{P}\left[-2\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)+2\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\Big\{2E_{P}\left[-\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)/2\right]+\lambda E_{P}\left[g(X)^{2}\right]\Big\}$
$=\operatorname*{arg\,min}_{\pi\in\{\pi(\dot{)}=\big(g(\dot{)}+1\big)/2\colon g\in{\mathcal{G}}\}}\Big\{4E_{P}\left[-\pi(X)Y_{1}-(1-\pi(X))Y_{0}\right]+\lambda E_{P}\left[\big(2\pi(X)-1\big)^{2}\right]\Big\}$
$=\widetilde{\pi}(\lambda).$
By defining ${\mathcal{G}}={\mathcal{G}}_{\Pi}$, the proof completes. ∎
@@
$\widehat{\pi}(\lambda)\coloneqq\operatorname*{arg\,max}_{\pi\in\Pi}\left\{\widehat{W}_{n}^{\mathrm{IPW}}(\pi)-\lambda\frac{1}{n}\sum^{n}_{i=1}\Big(\big(2\pi(X_{i})-1\big)^{2}\Big)\right\}.$
@@
$\widehat{g}(\lambda)=2\widehat{\pi}(4\lambda)-1.$
@@
which also implies $\widehat{g}(\lambda)\to 2\widehat{\pi}-1$ as $\lambda\to\infty$.
5 Computation: Relaxation of the NP-hard Problem
Maximizing $\widehat{W}(\pi)$ over $\pi\in\Pi$ is, in general, a combinatorial problem (e.g., empirical set selection), which is NP-hard for many natural classes $\Pi$. The equivalence result shows that one can instead solve least squares on a pseudo-outcome and then back out the policy.
6 Discussion
6.1 Conceptual unification.
EWM and plug-in policies are often portrayed as distinct. From the decision-making viewpoint, the former optimizes the ultimate goal directly, while the latter estimates the treatment effect as an intermediate quantity. Our results show that EWM can be written as least squares on a pseudo-outcome, with the policy embedded as $g=2\pi-1$. The difference is thus only an implementation detail. We conjecture that the gap in the theoretical results can also be further closed.
6.2 Beyond binary treatments.
For multiple treatments, we can derive corresponding equivalence results using one-vs-all pseudo-outcomes or a simplex-valued $g$ with squared loss on a vector CATE. The same idea extends to budget or fairness constraints by adding convex penalties to the least-squares objective.
@@ 7 Conclusion
We have shown that empirical welfare maximization and plug-in policy learning are equivalent after a simple reparameterization: EWM is least squares on a CATE target with predictors tied to the policy class. This equivalence clarifies theory, transfers regression guarantees to EWM, and yields a practical, convex training pipeline that avoids NP-hard search. We hope this unification simplifies both the analysis and implementation of policy learning in applied work.

### v2

@@ Abstract
The goal of policy learning is to train a policy function that recommends a treatment given covariates to maximize population welfare. There are two major approaches in policy learning: the empirical welfare maximization (EWM) approach and the plug-in approach. The EWM approach is analogous to a classification problem, where one first builds an estimator of the population welfare, which is a functional of policy functions, and then trains a policy by maximizing the estimated welfare. In contrast, the plug-in approach is based on regression, where one first estimates the conditional average treatment effect (CATE) and then recommends the treatment with the highest estimated outcome. This study bridges the gap between the two approaches by showing that both are based on essentially the same optimization problem. In particular, we prove an exact equivalence between EWM and least squares over a reparameterization of the policy class. As a consequence, the two approaches are interchangeable in several respects and share the same theoretical guarantees under common conditions. Leveraging this equivalence, we propose a regularization method for policy learning. The reduction to least squares yields a smooth surrogate that is typically easier to optimize in practice. At the same time, for many natural policy classes the inherent combinatorial hardness of exact EWM generally remains, so the reduction should be viewed as an optimization aid rather than a universal bypass of NP-hardness.
@@ 1 Introduction
    Equivalence. We establish that EWM over a policy class $\Pi$ is equivalent to least squares over the class ${\mathcal{G}}_{\Pi}\coloneqq\{g=2\pi-1:\pi\in\Pi\}$ for the target $Y_{1}-Y_{0}$. At the empirical level, EWM with an inverse-probability-weighted welfare estimator equals least squares of an IPW pseudo-outcome on $g\in{\mathcal{G}}_{\Pi}$, if the propensity score is known.
@@
    Computation. The equivalence provides a least–squares training objective whose continuous form enables standard optimization tools and simplifies implementation. For common discrete policy classes (e.g., trees, rule lists, set indicators), this surrogate often eases training in practice, but it does not in itself remove the intrinsic combinatorial hardness of exact EWM.
@@ 2 Setup
Let $e(x)=P(D=1\mid X=x)$ be the propensity score and $m_{d}(x)=E[Y_{d}\mid X=x]$ the conditional mean outcomes. We assume that there exists a constant $0<\epsilon<1/2$ independent of $n$ such that $\epsilon<e(X)<1-\epsilon$ almost surely and that the variables $X$ and $Y$ are bounded.
@@ 3.1 EWM Policy
$\widehat{W}_{n}^{\mathrm{IPW}}(\pi)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left(\frac{Y_{i}D_{i}}{e(X_{i})}\pi(X_{i})+\frac{Y_{i}(1-D_{i})}{1-e(X_{i})}\big(1-\pi(X_{i})\big)\right).$
@@
$\widehat{W}_{n}^{\mathrm{DR}}(\pi)\coloneqq$
$\frac{1}{n}\sum_{i=1}^{n}\left(\pi(X_{i})\left(\frac{D_{i}}{\widehat{e}(X_{i})}\big(Y_{i}-\widehat{m}_{1}(X_{i})\big)+\widehat{m}_{1}(X_{i})\right)+(1-\pi(X_{i}))\left(\frac{1-D_{i}}{1-\widehat{e}(X_{i})}\big(Y_{i}-\widehat{m}_{0}(X_{i})\big)+\widehat{m}_{0}(X_{i})\right)\right),$
where $\widehat{e}(x)$ is an estimator of the propensity score $e(x)$, and $\widehat{m}_{d}(x)$ ($d\in\{1,0\}$) is an estimator of the conditional expected outcome $m_{d}(x)$ (Athey & Wager, 2021). If the estimators of the nuisance parameters satisfy the Donsker condition or are constructed via sample splitting (Klaassen, 1987), also called cross-fitting (Chernozhukov et al., 2018), and meet mild convergence rate conditions, then the bias from estimation error vanishes at a faster rate than $1/\sqrt{n}$. We can estimate the bias-correction weight by using Riesz regression or density-ratio estimation, as proposed in Chernozhukov et al. (2024, 2022); Kato (2025a, b).
@@ 3.2 Plug-in Policy
$\widehat{m}_{d}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=d]\big(Y_{i}-f(X_{i})\big)^{2},$
where $\mathcal{F}$} is a class of regression functions $f\colon{\mathcal{X}}\to{\mathbb{R}}$, and construct $\widehat{\tau}$ as $\widehat{\tau}=\widehat{m}_{1}-\widehat{m}_{0}$. As another example, we can estimate $\tau(X)$ directly by regressing a variable that is (asymptotically) unbiased for $Y_{1,i}-Y_{0,i}$ on covariates $X_{i}$. For instance, using an IPW estimator for the target variable, we estimate $\tau$ as
$\widehat{\tau}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\sum^{n}_{i=1}\left(\frac{Y_{i}D_{i}}{e(X_{i})}-\frac{Y_{i}(1-D_{i})}{1-e(X_{i})}-f(X_{i})\right)^{2},$
@@ 4 Equivalence between EWM and least squares
For ${\mathcal{G}}_{\Pi}$, we have
$g^{*}=$ $\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[\Big(\big(Y_{1}-Y_{0}\big)-g(X)\Big)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[\big(Y_{1}-Y_{0}\big)^{2}-2g(X)\big(Y_{1}-Y_{0}\big)+g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+1\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)\right].$
@@
Continuing, we have
$g^{*}=$ $\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)+2\big(Y_{1}-Y_{0}\big)\right].$
We added and subtracted terms that are irrelevant to the optimization. Here, recall that
$\pi^{*}$ $=\operatorname*{arg\,min}_{\pi\in\left\{\pi(\cdot)=\big(g(\cdot)+1\big)/2\colon g\in{\mathcal{G}}_{\Pi}\right\}}E_{P}\left[-\pi(X)Y_{1}-(1-\pi(X))Y_{0}\right]$
$=\operatorname*{arg\,min}_{\pi\in\left\{\pi(\cdot)=\big(g(\cdot)+1\big)/2\colon g\in{\mathcal{G}}_{\Pi}\right\}}E_{P}\left[-\pi(X)(Y_{1}-Y_{0})-Y_{0}\right].$
Therefore, $g^{*}=2\pi^{*}-1$ holds, and the proof is complete. ∎
@@
We can similarly show the empirical version of this equivalence. Recall that given true $e(x)$, the EWM approach trains a policy $\widehat{\pi}_{n}$ as
@@ 4.2 $[0,1]$-policy
We next consider the case where $\pi$ is a $[0,1]$-valued function; that is, $\pi\colon{\mathcal{X}}\to[0,1]$. Let $\Pi$ be a policy class that contains such policies, $\pi\colon{\mathcal{X}}\to[0,1]$.
@@
We consider the equivalence between the following two problems: regularized EWM and modified least squares. We define the regularized EWM as follows:
@@
$\widetilde{g}(\lambda)=2\widetilde{\pi}(\lambda/4)-1.$
@@
which also implies $\widetilde{g}(\lambda)\to 2\pi^{*}-1$ as $\lambda\to 0$.
@@
For ${\mathcal{G}}_{\Pi}$, we have
$\widetilde{g}(\lambda)$ $=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[\left(\frac{1}{\sqrt{\lambda}}\big(Y_{1}-Y_{0}\big)-\sqrt{\lambda}g(X)\right)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[\frac{1}{\lambda}\big(Y_{1}-Y_{0}\big)^{2}-2g(X)\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2g(X)\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}E_{P}\left[-2\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)+2\big(Y_{1}-Y_{0}\big)+\lambda g(X)^{2}\right]$
$=\operatorname*{arg\,min}_{g\in{\mathcal{G}}_{\Pi}}\Big\{E_{P}\left[-\big(g(X)+1\big)\big(Y_{1}-Y_{0}\big)/2\right]+\lambda E_{P}\left[g(X)^{2}\right]/4\Big\}.$
We also have
$\widetilde{\pi}(\lambda)=\operatorname*{arg\,min}_{\pi\in\left\{\pi(\cdot)=\big(g(\cdot)+1\big)/2\colon g\in{\mathcal{G}}_{\Pi}\right\}}\Big\{E_{P}\left[-\pi(X)Y_{1}-(1-\pi(X))Y_{0}\right]+\lambda E_{P}\left[\big(2\pi(X)-1\big)^{2}\right]\Big\}.$
Since $g=2\pi-1$, the proof completes. ∎
@@
$\widehat{\pi}(\lambda)\coloneqq\operatorname*{arg\,max}_{\pi\in\Pi}\left\{\widehat{W}_{n}^{\mathrm{IPW}}(\pi)-\lambda\frac{1}{n}\sum^{n}_{i=1}\big(2\pi(X_{i})-1\big)^{2}\right\}.$
@@
$\widehat{g}(\lambda)=2\widehat{\pi}(\lambda/4)-1.$
@@
which also implies $\widehat{g}(\lambda)\to 2\widehat{\pi}-1$ as $\lambda\to 0$.
5 Computational implications
Maximizing $\widehat{W}(\pi)$ over $\pi\in\Pi$ is, in general, a combinatorial problem (e.g., empirical set selection) and is NP–hard for many natural classes $\Pi$. The equivalence established above replaces this discrete search with empirical risk minimization under a squared loss on a pseudo–outcome, followed by the simple back–mapping $g\mapsto\pi=(g+1)/2$. This yields a training pipeline that is substantially easier to optimize than direct 011 search.
The learning problem reduces to least squares regression of a pseudo–outcome, such as an IPW estimator, on a predictor $g\in\mathcal{G}_{\Pi}=\{2\pi-1:\pi\in\Pi\}$. This reformulation enables the use of standard continuous optimization tools (closed–form solvers, SGD, or second–order methods).
If the model class for $g$ is convex, then the least–squares is a convex optimization problem. In these cases, global optima are attained by off–the–shelf solvers. When $\Pi$ is discrete/nonconvex (e.g., trees, rule lists, or set indicators), the induced class $\mathcal{G}_{\Pi}$ is also nonconvex, so the objective need not be convex. Even then, the squared–loss formulation remains markedly easier to handle than maximizing a combinatorial welfare objective, because it replaces 011 structure with a smooth surrogate amenable to gradient–based methods.
Regardless of the specific function class, the least–squares view avoids the NP–hard policy search that arises in direct EWM and provides a stable, scalable ERM workflow with standard regularization and tuning.
@@ 6 Conclusion
From the decision-making, EWM and plug-in policies are often presented as distinct. From a decision-making perspective, the former optimizes welfare directly, whereas the latter bypasses the estimation of treatment effects as an intermediate step. We showed that this distinction collapses under a simple reparameterization: maximizing estimated welfare over a policy class is equivalent to least squares on a CATE target with predictors constrained to $g=2\pi-1$. This equivalence clarifies the relationship between the two traditions, allows regression guarantees to inform policy learning, and replaces combinatorial search with an empirical risk minimization routine that is typically easier to optimize in practice, while the underlying combinatorial hardness of exact policy search generally remains.
The same viewpoint extends beyond binary treatments. For multiple treatments, one can work with one-vs-all pseudo-outcomes or a simplex-valued predictor and a squared-loss on a vector CATE. Budget or fairness constraints can be incorporated through penalties in the least-squares objective. We hope this unification simplifies both analysis and implementation of policy learning in applied work, and we anticipate that remaining gaps in theoretical results can be narrowed under common regularity conditions.